# MLVTG: Mamba-Based Feature Alignment and LLM-Driven Purification for Multi-Modal Video Temporal Grounding

### v1

@@ 5 Ablation Study
R1 R5 mIoU mAP R1 mIoU mAP $\geq$ Very Good Avg.
@0.7 @0.7 @0.75 Avg. @0.7 @0.75 Avg HIT@1 mAP
[…]
@@
Figure 3: Effectiveness of LLM Pretrained Parameters.” means no LLM layer. “Rand” means randomly initialized.
LLM Param Frozen R1 R5 mAP
@0.7 @0.7 @0.75 @Avg.
43.5 59.4 39.0 37.7
Rand No 43.6 57.4 37.0 36.7
Rand Yes 43.3 57.9 38.1 37.1
Yes No 42.8 57.2 36.9 36.4
Yes Yes 45.1 61.4 39.3 38.4
Figure 4: Effectiveness of LLM Different Layer.
To validate the effectiveness of frozen pre-trained LLM parameters, we conduct experiments on the baseline with different initialization and freezing strategies. All models adopt a single-layer Transformer for fair comparison (Table 5). The baseline achieves 38.4 mAP@Avg. Adding a trainable random initialized layer reduces performance by 2.7% (row 2 vs. 1), showing that structural imitation introduces noise. Freezing random initialized layers yields 37.1 (–3.4%, row 3 vs. 5), highlighting the need for pre-trained LLM parameters. Full fine-tuning leads to 36.4 (–5.2%, row 4 vs. 5), confirming that updates disrupt alignment. In contrast, frozen LLMs achieve 45.1 R1@0.7, 38.4 mAP@Avg, and 61.4 R5@0.7, demonstrating better temporal grounding.
Layer-wise analysis reveals that the multi-modal alignment ability of LLMs is highly dependent on layer selection. As shown in Fig 4, using the 20th layer’s frozen parameters yields the best performance. In contrast, higher layer (e.g., 25th) lead to significant degradation, suggesting that excessive abstraction harms visual grounding. Overall, performance shows a parabolic trend across layers, with the range 15–22th forming a “semantic sweet spot” that avoids both syntactic noise from lower layers and semantic oversimplification from higher ones.
@@ 6 Visualization
Figure 5: Feature Alignment Visualization. Darker colors indicate lower similarity.
Figure 5 illustrates the cosine similarity heatmap between query and video features. Initially, the alignment between these features is poor, as evidenced by the scattered distribution of similarity scores. Upon introducing MambaAligner, a substantial improvement in the alignment quality between video and textual features is observed. Nevertheless, the presence of spurious regions exhibiting high similarity, particularly in the early video clip due to background distractors, reveals limitations in alignment precision. Subsequent refinement by LLMRefiner markedly enhances alignment accuracy by effectively suppressing noise and concentrating high-similarity regions within the ground-truth interval (63–132 seconds). The visualization explicitly demonstrates the incremental benefit provided by MambaAligner for enhancing alignment and the critical refinement contributed by LLMRefiner for precise localization. Furthermore, Fig. 6 further compares our method MLVTG with the baseline on QVHighlights.
Figure 6: Visualization of joint temporal localization and highlight detection on QVHighlights.
@@ 7 Conclusion
This paper proposes MLVTG, a novel framework for video temporal grounding that combines Mamba and frozen LLM layers for effective multi-modal alignment. Experiments on three benchmarks show that MLVTG outperforms state-of-the-art methods in both temporal localization and highlight detection. Ablation studies further confirm the effectiveness of each component, demonstrating the robustness and superiority of MLVTG. Future work will explore extending MLVTG to incorporate audio modalities for more comprehensive video temporal grounding.

### v2

@@ 3 Methodology
R1 mIoU mAP R1 mIoU mAP $\geq$ Very Good Avg.
@0.7 Avg. @0.7 Avg. HIT@1 mAP
43.5 54.7 39.0 35.6 49.4 39.0 64.2 39.3 72.7
48.9 59.1 42.4 38.7 50.2 40.4 64.8 40.1 77.1
45.1 55.8 39.3 35.4 49.4 39.4 62.8 38.2 77.7
50.5 59.9 44.6 38.7 50.3 40.8 65.2 40.5 80.1
@@
Fig. 7: Effectiveness of LLM Pretrained Parameters.” means no LLM layer. “Rand” means randomly initialized.
LLM Param Frozen R1 mAP
@0.7 @0.75 @Avg.
43.5 39.0 37.7
Rand No 43.6 37.0 36.7
Rand Yes 43.3 38.1 37.1
Yes No 42.8 36.9 36.4
Yes Yes 45.1 39.3 38.4
3.3 LLMRefiner.
To refine the semantics of latent representations and enhance multi-modal alignment, we propose the Mamba-based LLMRefiner module. This design addresses the performance deviations caused by structural differences. The module integrates two linear layers $F_{L}^{1}$ and $F_{L}^{2}$, and a pre-trained Mamba-based large language model layer $F_{LLM}$ to improve semantic representation and multi-modal alignment. Leveraging the inherent strong multi-modal reasoning capabilities of large language models, the frozen-parameter $F_{LLM}$ layer effectively transfers textual priors into the visual domain, filtering out redundant or irrelevant information in video content.
$\mathbf{Z}_{refine}=F_{L}^{1}*F_{LLM}(\mathbf{Z}_{out})*F_{L}^{2}.$ (6)
Fig. 8: Efficiency Analysis of MambaAligner.
Fig. 9: Effectiveness of LLM Different Layers.
Fig. 10: Feature Alignment Visualization. Darker colors indicate lower similarity.
3.4 Joint Optimization for TL & HD.
$\mathcal{L}_{overall}$ $=\lambda_{f}\mathcal{L}_{f}+\lambda_{reg}\mathcal{L}_{reg}+\lambda_{1}\mathcal{L}^{inter}+\lambda_{2}\mathcal{L}^{intra},$ (7)
where $\mathcal{L}_{f}$ represents the classification loss, and $\mathcal{L}_{reg}$ denotes the regression loss, both dedicated to optimizing TL. Conversely, $\mathcal{L}^{inter}$ and $\mathcal{L}^{intra}$ are explicitly employed to enhance the accuracy and discriminative power of HD [18].
4 Experiments
4.1 Implementation Details.
We use CLIP and SlowFast features from pre-segmented video clips at fixed frame rates (0.5/1 FPS), while text features are extracted through the CLIP text encoder. The model employs a MambaAligner (4 blocks, hidden size 1024) and an LLMRefiner which operates on 2056 dimension features extracted from the 20th layer. Training used a batch size of 32 for 200 epochs on a RTX4090 GPU.
4.2 Comparison with Start-of-the-Arts
Table 5 demonstrates outstanding performance on QVHighlights. On the test set, it achieves the highest mAP@0.75 and HD mAP. Since TR-DETR specifically designs a HD module to assist temporal localization, it performs well. However, more generalized MLVTG outperforms it on the generic TL dataset Charades-STA, leading QD-DETR by 8.6% in R1@0.7 and outperforming both UVCOM and QD-DETR in HD task HIT@1. On the Charades-STA, MLVTG achieves optimal results under R1@0.7 and mIoU in Table 5. Moreover, without relying on complex detection architectures, it delivers superior overall performance compared to detector-based methods such as TR-DETR. On TVSum , MLVTG attains an average AP of 80.1,matching SOTA CO-AV in Table 5. Even without multi-dataset training like UniVTG, MLVTG remains highly competitive. And without introducing audio modality, it exhibits stronger generalization capability than complex models such as Joint-VA.
5 Ablation Study
As shown in the Table 7, MambaAligner alone significantly improves temporal localization performance: on QVHighlights, R1@0.7 increased by 12.4%, mIoU by 8.0%, and mAP@Avg by 8.7%; on Charades-STA, it improved by 8.7%, 1.6%, and 3.6%, respectively. LLMRefiner provides improvements (e.g., +3.7% in R1@0.7 on QVHighlights), and the combination of both yields the best results, with total gains of 16.1% in R1@0.7, 9.5% in mIoU, and 14.4% in mAP@Avg. On the HD task, the full model also achieves the highest performance (+3.1% in mAP and +10.2% in Avg). Furthermore, Table 7 demonstrate that using frozen pre-trained LLM parameters delivers the best outcomes (e.g., 45.1 R1@0.7 and 38.4 mAP@Avg), while fine-tuning or random initialization leads to degradation.
Fig 8 presents that MambaAligner demonstrates significantly lower VRAM usage and inference time compared to the baseline with longer sequence length. Although current video datasets consist of relatively short sequences, which limits full demonstration of its linear complexity advantage, the design exhibits strong potential for high efficiency in long-sequence processing. Fig 9 further reveals that using the frozen parameters from the 20th layer performs optimally, with layers 15–22 forming a "semantic sweet spot"; either higher or lower layers result in decreased performance.
@@ 6 Visualization
As shown in Figure 10, the cosine similarity heatmap between query and video features reveals poor initial alignment. After introducing MambaAligner, a substantial improvement in alignment quality is observed, though spurious high-similarity regions remain—particularly in the early clips due to background distractors. Further refinement with LLMRefiner effectively suppresses noise and concentrates high similarity within the ground-truth interval (63–132 seconds), resulting in more precise alignment.
@@ 7 Conclusion
This paper proposes MLVTG, a novel framework for video temporal grounding that presents MambaAligner and LLMRefiner for effective multi-modal alignment. Experiments on three benchmarks show that MLVTG outperforms highly competitive performance in both temporal localization and highlight detection. Ablation studies further confirm the effectiveness of each component, demonstrating the robustness and superiority of MLVTG. Future work will explore extending MLVTG to incorporate audio modalities for more comprehensive video temporal grounding.